# Neural Topic Modeling with large language models in the loop

### v1

@@ 5.1 Experimental Setup
We conduct experiments on four widely used datasets in topic modeling, including 20Newsgroup (Lang, 1995) (20News), Reuters-21578 (Aletras & Stevenson, 2013) (R8), DBpedia (Auer et al., 2007) and AGNews (Zhang et al., 2015). Further details of these datasets are described in the Appendix D.1. The number of mined topics (i.e., $K$) is commonly regarded as a hyper-parameter for the dataset (Zhao et al., 2020; Wu et al., 2024). For datasets containing long documents, such as 20News and R8, we set the number of topics to 50. For datasets with short documents, such as DBpedia and AGNews, we set the number to 25. We also run experiments at different $K$ values, which are reported in Appendix E.
@@
We compare LLM-ITL with topic models of different types, including Latent Dirichlet Allocation (LDA) (Blei et al., 2003); Neural Variational Document Model (NVDM) (Miao et al., 2017); LDA with Products of Experts (PLDA) (Srivastava & Sutton, 2017); Embedded Topic Model (ETM) (Dieng et al., 2020); Neural Topic Model with Covariates, Supervision, and Sparsity (SCHOLAR) (Card et al., 2017); Contrastive Learning Neural Topic Model (CLNTM) (Nguyen & Luu, 2021); BERTopic (Grootendorst, 2022) and TopicGPT (Pham et al., 2023). Further details about these baselines and their settings are provided in Appendix D.2.
@@
LLM-ITL is a framework compatible with most NTMs and LLMs. We use ETM, SCHOLAR, and CLNTM as the base models for our experiments. We use LLAMA3-8B-Instruct [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct in LLM-ITL for main experiments. For OT computation, we use GloVe (Pennington et al., 2014) word embeddings pre-trained on Wikipedia to construct the OT cost matrix, and compute the OT distance using the POT [^3] https://pythonot.github.io/ package. For the topic labeling confidence, we use label token probability in Eq. 8 for our main experiments. As for the hyper-parameters of LLM-ITL, we set the topic refinement strength $\gamma$ to 200; and the refinement step $T^{\text{refine}}$ to 150 for ETM and 450 for SCHOLAR and CLNTM. We set the number of words for the topic label to 2 when prompting the LLM. All hyper-parameters of LLM-ITL are studied in the following sections. As for the LLM generation, we use greedy decoding to enable deterministic output and set the maximum new generation tokens to 300. Each trial [^4] All experiments are conducted five times with different model random seeds throughout this work. The mean and standard deviation values of performance are reported. of LLM-ITL in our experiment takes a few hours on a single 80GB A100 GPU.
@@
We evaluate both the topic quality and the document representation quality for topic models. For topic quality, we apply the widely used topic coherence metric, Normalized Pointwise Mutual Information (NPMI) (Lau et al., 2014). We report the average NPMI values (in percentage) of all topics. Moreover, topic diversity (Dieng et al., 2020) is also evaluated, which is reported in Appendix E.3. For document representation quality, we evaluate the alignment between a document’s true label and the top-weighted topic of its topical representation using external clustering metrics, known as topic alignment (Chuang et al., 2013; Pham et al., 2023). We compute the commonly used Purity and Normalized Mutual Information (NMI) to evaluate clustering performance. Since Purity and NMI are considered equally important, fall within the same range (from 0 to 1), and are often reported together, we report their average as PN, serving as an overall indicator of topic alignment performance. Detailed results for Purity and NMI are provided in Appendix E.4. Intuitively, topic coherence (e.g., NPMI) reflects how coherent the learned topic words are, while topic alignment (e.g., PN) indicates how well the model represents the documents through the learned topics. Further details on the calculation of these metrics are provided in Appendix D.3.
@@ 5.2 Results
Table 1: Topic coherence (NPMI) and topic alignment (PN). The best and second-best performance of each column are highlighted in boldface and underlined, respectively. “NA” indicates the evaluation is not applicable. The performance improvement of LLM-ITL over its base model is computed.
@@
NPMI PN NPMI PN NPMI PN NPMI PN
LDA (Blei et al., 2003) 3.95 ± 0.27 0.489 ± 0.009 -3.43 ± 0.49 0.700 ± 0.005 6.42 ± 0.22 0.762 ± 0.013 7.74 ± 0.44 0.596 ± 0.008
NVDM (Miao et al., 2017) -16.89 ± 0.71 0.145 ± 0.006 -8.12 ± 0.62 0.360 ± 0.012 -7.82 ± 0.49 0.185 ± 0.004 -7.39 ± 0.85 0.248 ± 0.013
PLDA (Srivastava & Sutton, 2017) -13.70 ± 0.76 0.101 ± 0.005 -6.26 ± 0.45 0.524 ± 0.009 5.60 ± 0.58 0.653 ± 0.008 7.03 ± 0.77 0.487 ± 0.011
BERTopic (Grootendorst, 2022) 2.18 ± 0.73 0.342 ± 0.008 -0.65 ± 0.14 0.697 ± 0.001 7.68 ± 0.84 0.720 ± 0.009 5.00 ± 0.92 0.450 ± 0.010
@@
ETM (Dieng et al., 2020) 2.96 ± 0.42 0.404 ± 0.010 -1.71 ± 0.72 0.669 ± 0.009 4.49 ± 0.45 0.762 ± 0.011 6.15 ± 0.59 0.568 ± 0.008
LLM-ITL (ETM) 8.92 ± 0.74 0.398 ± 0.010 7.13 ± 0.57 0.686 ± 0.012 14.83 ± 0.79 0.742 ± 0.016 12.04 ± 0.95 0.569 ± 0.005
$\uparrow$ 5.96 $\downarrow$ 0.006 $\uparrow$ 8.84 $\uparrow$ 0.017 $\uparrow$ 10.34 $\downarrow$ 0.020 $\uparrow$ 5.89 $\uparrow$ 0.001
SCHOLAR (Card et al., 2017) -2.12 ± 0.77 0.582 ± 0.010 -4.06 ± 0.15 0.680 ± 0.013 12.32 ± 1.54 0.825 ± 0.015 6.41 ± 0.70 0.638 ± 0.003
LLM-ITL (SCHOLAR) 7.58 ± 0.45 0.568 ± 0.010 -0.78 ± 0.60 0.680 ± 0.012 15.13 ± 1.61 0.828 ± 0.013 11.07 ± 0.78 0.639 ± 0.002
$\uparrow$ 9.70 $\downarrow$ 0.014 $\uparrow$ 3.28 $\uparrow$ $0$uparrow 2.81 $\uparrow$ 0.003 $\uparrow$ 4.66 $\uparrow$ 0.001
CLNTM (Nguyen & Luu, 2021) -2.21 ± 1.07 0.575 ± 0.011 -4.99 ± 0.36 0.691 ± 0.005 3.75 ± 1.35 0.683 ± 0.040 5.20 ± 1.38 0.607 ± 0.014
LLM-ITL (CLNTM) 8.12 ± 0.49 0.576 ± 0.005 -1.25 ± 0.57 0.691 ± 0.005 10.03 ± 1.11 0.684 ± 0.039 11.13 ± 0.96 0.594 ± 0.011
$\uparrow$ 10.33 $\uparrow$ 0.001 $\uparrow$ 3.74 $\uparrow$ $0$uparrow 6.28 $\uparrow$ 0.001 $\uparrow$ 5.93 $\downarrow$ 0.013
@@
We show the performance of topic coherence and topic alignment for different models in Table 1. We summarize the following remarks based on the results: (1) As for topic coherence, LLM-ITL significantly improves the performance of the base model and achieves state-of-the-art (SOTA) performance. (2) In terms of topic alignment, LLM-ITL inherits the document representation capability of its base model and shows SOTA performance in most cases. (3) Moreover, for long-document corpora such as 20News and R8, LLM-ITL outperforms existing LLM-based topic models like TopicGPT in terms of topic alignment, where the LLM alone may fail to fully capture the topics for long documents. Additional results about topic coherence (Appendix E.1) and alignment (Appendix E.2) at different settings of $K$, as well as topic diversity (Appendix E.3) performance, are illustrated in the appendix.
Figure 2: Learning curves of LLM-ITL with different base models in terms of topic coherence (NPMI) and topic alignment (PN) on 20News (Top) and DBpedia (Bottom). The gray and brown curves indicate the status of the base model and LLM-ITL, respectively. The base models are (a) ETM, (b) SCHOLAR and (c) CLNTM, respectively.
Learning Status of LLM-ITL
To clearly demonstrate how LLM-ITL improves its base model, we illustrate the learning curves for topic coherence and alignment on both a long document dataset (e.g., 20News) and a short document dataset (e.g., DBpedia) in Figure 2. We have the following observations based on the results: (1) When topic refinement is applied, LLM-ITL significantly improves the topic coherence  [^5] Topic coherence is not correlated with the training of topic models (Chang et al., 2009), which can result in a drop in coherence in the learning curve. of the base model. (2) LLM-ITL has little overall influence on topic alignment in most cases.
Figure 3: Learning curves of LLM-ITL (ETM) with different $T^{\text{refine}}$ in terms of topic coherence (NPMI) and topic alignment (PN) on 20News ((a) and (b)) and DBpedia ((c) and (d)).
@@
As indicated by previous works (Bhatia et al., 2017; Yang et al., 2024), a topic model with better topic coherence may not perform well in document representations (i.e., topic alignment) at the same time, and vice versa. To provide further insights into how LLM-ITL balances between topic coherence and alignment, we illustrate the learning curves of LLM-ITL (ETM) with different $T^{\text{refine}}$ in terms of both metrics. As illustrated in Figure 3, we have the following observations: (1) While starting topic refinement earlier (e.g., $T^{\text{refine}}=5$) can lead to greater improvements in topic coherence, it may also introduce more irrelevant information about the corpus that is from the LLM’s knowledge, thereby harming topic alignment performance. (2) For larger values of $T^{\text{refine}}$, the performance in terms of both topic coherence and alignment is comparable, indicating little sensitivity to the settings of $T^{\text{refine}}$ in a certain range. These observations suggest the effectiveness of $T^{\text{refine}}$ in balancing topic coherence and alignment in LLM-ITL. For more hyper-parameter studies of LLM-ITL, see Appendix G.
Table 2: Examples of different topic models’ output for a given document from 20News. Only the document’s top assigned/weighted ($>=0.1$) topics of its topical proportion/representation are listed in LDA and LLM-ITL.
@@
During the inference phase, LLM-ITL infers the topic proportion (i.e., topical representation) for a given document from the NTM component, and obtains the topic label from the LLM component, as shown in Table 5.2 (where ETM is used as the base model). We can observe that (1) Compared to topic models with top-words topics such as LDA, LLM-ITL provides more coherent topic words and offers topic labels, making the semantic meaning of the topics easier to identify. (2) Compared to the LLM-based topic model TopicGPT, LLM-ITL can obtain topic proportions as an indicator of the importance or relevance of topics to the document, offering more practical usage. For example, for TopicGPT, “Internet Culture” should be less relevant for the example document than “Software Development” if a good topic proportion is available.
Figure 4: Learning curves of LLM-ITL (ETM) with different LLMs in terms of topic coherence (NPMI) and topic alignment (PN) on 20News ((a) and (b)) and DBpedia ((c) and (d)).
@@
LLM-ITL is a framework compatible with most LLMs. Here, we examine the flexibility of LLM-ITL by integrating it with various LLMs. Apart from LLAMA3-8B-Instruct [^6] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct (Dubey et al., 2024), we implement LLM-ITL with the latest open-sourced LLMs, including Mistral-7B-Instruct-v0.3777https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), Phi-3-Mini-128K-Instruct [^8] https://huggingface.co/microsoft/Phi-3-mini-128k-instruct (Abdin et al., 2024), Yi-1.5-9B-Chat [^9] https://huggingface.co/01-ai/Yi-1.5-9B-Chat (Young et al., 2024), Qwen1.5-32B-Chat [^10] https://huggingface.co/Qwen/Qwen1.5-32B-Chat (Bai et al., 2023) and LLAMA3-70B-Instruct [^11] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct (Dubey et al., 2024). As shown in Figure 4, LLM-ITL consistently improves topic coherence of its base model across different LLMs, and the improvement can be further enhanced when using larger LLMs such as LLAMA3-70B-Instruct, demonstrating the flexibility of LLM-ITL.
Figure 5: Ablation studies. First row: Ablation for confidence on 20News ((a) and (b)) and DBpedia ((c) and (d)); Second row: Ablation for OT on 20News ((e) and (f)) and DBpedia ((g) and (h)). Error bars are omitted for clarity in the figure.
@@
Here, we investigate the effectiveness of including the confidence scores during topic refinement. We apply No Conf. (i.e., $\text{Conf}(\bm{w}^{l})=1$ in Eq. 10 for all refinement), Label Token Prob. (i.e., Eq. 8) and Word Intrusion Conf. (i.e., Eq. 9) to LLM-ITL (ETM). We plot the learning curves for both metrics on 20News and DBpedia. From the results in Figure 5 (first row), we can observe that label token probability and word intrusion confidence consistently yield better performance in terms of PN. This suggests that by including proposed confidence during topic refinement, we reduce potential noisy suggestions from the LLM and achieve better topical representation for documents. For further studies on alternative LLM confidence measures, see Appendix F.
Ablation Study for OT
Here, we study the effectiveness of our OT-based topic refinement. We apply different metrics to measure the difference between the topic word distributions from the NTM and those from the LLM, including Kullback–Leibler (KL) divergence, Jensen–Shannon Divergence (JSD), Hellinger Distance (HD), and Total Variation Distance (TVD). As illustrated in Figure 5 (second row), our OT-based approach significantly benefits topic coherence compared to other distribution measurements.

### v2

@@ 5.1 Experimental Setup
We conduct experiments on four widely used and publicly available datasets for topic modeling, including 20Newsgroup (Lang, 1995) (20News), Reuters-21578 (Aletras and Stevenson, 2013) (R8), DBpedia (Auer et al., 2007) and AGNews (Zhang et al., 2015). Further details of these datasets are described in the Appendix D.1. The number of mined topics (i.e., $K$) is commonly regarded as a hyper-parameter for the dataset (Zhao et al., 2020; Wu et al., 2024a). For datasets containing long documents, such as 20News and R8, we set the number of topics to 50. For datasets with short documents, such as DBpedia and AGNews, we set the number to 25. We also run experiments at different $K$ values, which are reported in Appendix E.
@@
The LLM-ITL framework is highly modular and can be seamlessly integrated with a wide range of NTMs. We integrated LLM-ITL with 8 commonly benchmarked NTMs, as listed in Table 1. We also evaluate other types of topic models, including probabilistic models such as LDA (Blei et al., 2003), cluster-based models like BERTopic (Grootendorst, 2022), and the latest LLM-based model, TopicGPT (Pham et al., 2023). Further details about these baselines and their settings are provided in Appendix D.2.
@@
We use LLAMA3-8B-Instruct (Dubey et al., 2024) in LLM-ITL for main experiments. For LLM generation, we use greedy decoding to enable deterministic output and set the maximum new generation tokens to 300. For OT computation, we use GloVe (Pennington et al., 2014) word embeddings pre-trained on Wikipedia to construct the OT cost matrix, and compute the OT distance using the POT (Flamary et al., 2021) package. For topic labeling confidence, we use label token probability in Eq. 8 for our main experiments. For hyper-parameters of LLM-ITL, we maintain consistent settings across all datasets and base models it integrates with: the topic refinement strength $\gamma$ is set to 200; the number of words for the topic label is set to 2 when prompting the LLM; the warm-up steps are set as $T^{\text{refine}}=T^{\text{total}}-50$, where $T^{\text{total}}$ denotes the total training steps of the base model without integration. Each component and hyper-parameter of LLM-ITL is studied in the following sections. Each trial [^2] All experiments are conducted five times with different model random seeds throughout this work. The mean and standard deviation values of performance are reported. of LLM-ITL in our experiment takes a few hours on a single 80GB A100 GPU.
@@
We evaluate both the topic quality and the document representation quality for topic models. For topic quality, we apply the widely used topic coherence metric, $\bm{C_{V}}$ (Röder et al., 2015). We report the average $C_{V}$ values of all learned topics. For document representation quality, we evaluate the alignment between a document’s true label and the top-weighted topic of its topical representation, known as topic alignment (Chuang et al., 2013; Pham et al., 2023). This is commonly evaluated using external clustering metrcis, Purity and Normalized Mutual Information (NMI). Since Purity and NMI are considered equally important, fall within the same range (from 0 to 1), and are often reported together, we report their average as PN, serving as an overall indicator of topic alignment performance. Detailed results for Purity and NMI are provided in Appendix E.5. Intuitively, topic coherence ($C_{V}$) reflects how coherent the learned topic words are, while topic alignment (PN) indicates how well the model represents the documents through the learned topics. Further details on the calculation of these metrics are provided in Appendix D.3. In addition, we compute the performance improvement of LLM-ITL over its base model as: $\frac{S(\text{base+LLM-ITL})-S(\text{base})}{S(\text{base})}$, where $S(\cdot)$ represents the evaluation metric. As for other topic model evaluation metrics, including topic diversity (TD) (Dieng et al., 2020) and overall topic quality (TQ) (Wang et al., 2022) are reported in Appendix E.3 and E.4, respectively.
@@ 5.2 Results
$\bm{C_{V}}$ PN $\bm{C_{V}}$ PN $\bm{C_{V}}$ PN $\bm{C_{V}}$ PN
LDA (Blei et al., 2003) 0.529 ± 0.006 0.489 ± 0.009 0.384 ± 0.006 0.700 ± 0.005 0.537 ± 0.010 0.762 ± 0.013 0.544 ± 0.013 0.596 ± 0.008
BERTopic (Grootendorst, 2022) 0.404 ± 0.005 0.342 ± 0.008 0.323 ± 0.010 0.697 ± 0.001 0.545 ± 0.015 0.720 ± 0.009 0.506 ± 0.016 0.450 ± 0.010
@@
NVDM (Miao et al., 2017) 0.261 ± 0.008 0.141 ± 0.006 0.298 ± 0.010 0.359 ± 0.010 0.329 ± 0.017 0.175 ± 0.009 0.355 ± 0.010 0.252 ± 0.009
+ LLM-ITL 0.336 ± 0.014 0.138 ± 0.007 0.374 ± 0.016 0.360 ± 0.011 0.479 ± 0.025 0.170 ± 0.006 0.455 ± 0.016 0.248 ± 0.013
$\uparrow$ 28.7% $\downarrow$ 2.1% $\uparrow$ 25.5% $\uparrow$ 0.3% $\uparrow$ 45.6% $\downarrow$ 2.9% $\uparrow$ 28.2% $\downarrow$ 1.6%
PLDA (Srivastava and Sutton, 2017) 0.368 ± 0.011 0.272 ± 0.008 0.375 ± 0.004 0.525 ± 0.006 0.502 ± 0.014 0.664 ± 0.009 0.551 ± 0.012 0.481 ± 0.010
+ LLM-ITL 0.525 ± 0.010 0.272 ± 0.011 0.466 ± 0.011 0.522 ± 0.008 0.627 ± 0.020 0.660 ± 0.005 0.625 ± 0.017 0.486 ± 0.008
$\uparrow$ 42.7% $\uparrow$ 0.0% $\uparrow$ 24.3% $\downarrow$ 0.6% $\uparrow$ 24.9% $\downarrow$ 0.6% $\uparrow$ 13.4% $\uparrow$ 1.0%
SCHOLAR (Card et al., 2017) 0.479 ± 0.019 0.582 ± 0.010 0.386 ± 0.004 0.680 ± 0.013 0.608 ± 0.019 0.825 ± 0.015 0.576 ± 0.011 0.638 ± 0.003
+ LLM-ITL 0.591 ± 0.012 0.568 ± 0.010 0.439 ± 0.008 0.680 ± 0.012 0.678 ± 0.017 0.828 ± 0.013 0.655 ± 0.012 0.639 ± 0.002
$\uparrow$ 23.4% $\downarrow$ 2.4% $\uparrow$ 13.7% $\uparrow$ 0.0% $\uparrow$ 11.5% $\uparrow$ 0.4% $\uparrow$ 13.7% $\uparrow$ 0.2%
ETM (Dieng et al., 2020) 0.491 ± 0.006 0.404 ± 0.010 0.433 ± 0.006 0.679 ± 0.018 0.513 ± 0.004 0.762 ± 0.011 0.534 ± 0.013 0.568 ± 0.008
+ LLM-ITL 0.578 ± 0.008 0.398 ± 0.010 0.571 ± 0.010 0.686 ± 0.012 0.704 ± 0.015 0.742 ± 0.016 0.644 ± 0.017 0.569 ± 0.005
$\uparrow$ 17.7% $\downarrow$ 1.5% $\uparrow$ 31.9% $\uparrow$ 1.0% $\uparrow$ 37.2% $\downarrow$ 2.6% $\uparrow$ 20.6% $\uparrow$ 0.2%
NSTM (Zhao et al., 2020) 0.444 ± 0.015 0.373 ± 0.005 0.412 ± 0.004 0.665 ± 0.014 0.652 ± 0.005 0.767 ± 0.011 0.588 ± 0.022 0.593 ± 0.010
+ LLM-ITL 0.521 ± 0.015 0.360 ± 0.006 0.549 ± 0.011 0.673 ± 0.006 0.700 ± 0.013 0.752 ± 0.014 0.675 ± 0.012 0.585 ± 0.010
$\uparrow$ 17.3% $\downarrow$ 3.5% $\uparrow$ 33.3% $\uparrow$ 1.2% $\uparrow$ 7.4% $\downarrow$ 2.0% $\uparrow$ 14.8% $\downarrow$ 1.3%
CLNTM (Nguyen and Luu, 2021) 0.490 ± 0.014 0.575 ± 0.011 0.361 ± 0.008 0.691 ± 0.005 0.500 ± 0.015 0.683 ± 0.040 0.558 ± 0.028 0.607 ± 0.014
+ LLM-ITL 0.612 ± 0.010 0.576 ± 0.005 0.436 ± 0.008 0.691 ± 0.005 0.612 ± 0.019 0.684 ± 0.039 0.655 ± 0.020 0.594 ± 0.011
$\uparrow$ 24.9% $\uparrow$ 0.2% $\uparrow$ 20.8% $\uparrow$ 0.0% $\uparrow$ 22.4% $\uparrow$ 0.1% $\uparrow$ 17.4% $\downarrow$ 2.1%
WeTe (Wang et al., 2022) 0.495 ± 0.017 0.318 ± 0.009 0.529 ± 0.011 0.649 ± 0.002 0.546 ± 0.003 0.721 ± 0.012 0.560 ± 0.006 0.516 ± 0.004
+ LLM-ITL 0.583 ± 0.028 0.316 ± 0.008 0.599 ± 0.015 0.656 ± 0.005 0.613 ± 0.015 0.703 ± 0.012 0.623 ± 0.007 0.517 ± 0.004
$\uparrow$ 17.8% $\downarrow$ 0.6% $\uparrow$ 13.2% $\uparrow$ 1.1% $\uparrow$ 12.3% $\downarrow$ 2.5% $\uparrow$ 11.2% $\uparrow$ 0.2%
ECRTM (Wu et al., 2023) 0.323 ± 0.014 0.516 ± 0.005 0.308 ± 0.006 0.673 ± 0.008 0.581 ± 0.026 0.668 ± 0.202 0.438 ± 0.035 0.554 ± 0.019
+ LLM-ITL 0.551 ± 0.025 0.519 ± 0.002 0.364 ± 0.010 0.667 ± 0.007 0.684 ± 0.013 0.696 ± 0.142 0.505 ± 0.025 0.553 ± 0.020
$\uparrow$ 70.6% $\uparrow$ 0.6% $\uparrow$ 18.2% $\downarrow$ 0.9% $\uparrow$ 17.7% $\uparrow$ 4.2% $\uparrow$ 15.3% $\downarrow$ 0.2%
Table 1: Topic coherence ($C_{V}$) and topic alignment (PN). “NA” indicates the evaluation is not applicable. The performance improvement of LLM-ITL over its base model is computed.
@@
We show the performance of topic coherence and topic alignment for different models in Table 1. We summarize the following remarks based on the results: (1) For topic coherence, using LLM-ITL significantly improves performance across all cases, with minimum gains of +7.4% (NSTM as the base model on DBpedia) and maximum gains of +70.6% (ECRTM as the base model on 20News) over the base model. (2) In terms of topic alignment, integrating LLM-ITL demonstrates comparable performance to its base model, with changes ranging from -3.5% to +4.2%. (3) Moreover, for long-document corpora such as 20News and R8, applying LLM-ITL outperforms existing LLM-based topic models like TopicGPT in terms of topic alignment in most cases. This suggests that LLMs alone struggle to fully capture the topics of long documents, highlighting the advantages of integrating NTMs to enhance topical representation for document collections. Additional results with different settings and evaluation metrics are presented in Appendix E.
Figure 2: Learning curves of LLM-ITL (ETM as the base model) with different $T^{\text{refine}}$ in terms of topic coherence ($C_{V}$) and topic alignment (PN) on 20News.
@@
Over-reliance on the LLM’s refinement may introduce out-of-corpus information, thereby harming topical representation of documents for the corpus (i.e., topic alignment). Here, we demonstrate the effectiveness of using warm-up integration to balance topic coherence and alignment. From the learning curves illustrated in Figure 2, we observe that starting topic refinement earlier (e.g., $T^{\text{refine}}=5$) can result in greater improvements in topic coherence but leads to poorer topic alignment performance. For larger values of $T^{\text{refine}}$, the performance in terms of both metrics becomes better balanced and comparable, indicating the effectiveness of warm-up integration and low sensitivity within a certain range. For more hyper-parameter studies of LLM-ITL, see Appendix G.
@@
Table 2: Examples of different topic models’ output for a given document from 20News. Only the document’s top assigned/weighted ($>=0.1$) topics of its topical proportion/representation are listed in LDA and LLM-ITL (with ETM as the base model).
@@
During the inference phase, LLM-ITL infers the topic proportion for a given document from the NTM component, and obtains the topic label from the LLM component, as shown in Table 5.2. We can observe that (1) Compared to topic models with top-words topics such as LDA, LLM-ITL provides more coherent topic words and offers topic labels, making the semantic meaning of the topics easier to identify. (2) Compared to the LLM-based topic model TopicGPT, LLM-ITL can obtain topic proportions as an indicator of the importance or relevance of topics to the document, offering more practical usage. For example, for TopicGPT, “Internet Culture” should be less relevant for the example document than “Software Development” if a good topic proportion is available.
Figure 3: Learning curves of LLM-ITL (ETM as the base model) with different LLMs in terms of topic coherence ($C_{V}$) and topic alignment (PN) on 20News.
@@
LLM-ITL is a framework compatible with most LLMs. Here, we examine the flexibility of LLM-ITL by integrating it with various LLMs. Apart from LLAMA3-8B-Instruct (Dubey et al., 2024), we implement LLM-ITL with the latest open-sourced LLMs, including Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), Phi-3-Mini-128K-Instruct (Abdin et al., 2024), Yi-1.5-9B-Chat (Young et al., 2024), Qwen1.5-32B-Chat (Bai et al., 2023) and LLAMA3-70B-Instruct (Dubey et al., 2024). As shown in Figure 3, LLM-ITL consistently improves topic coherence of its base model across different LLMs, and the improvement can be further enhanced when using larger LLMs such as LLAMA3-70B-Instruct.
Figure 4: Ablation for confidence. Error bars are omitted for clarity in the figure.
@@
Here, we investigate the effectiveness of including the confidence scores during topic refinement. As illustrated in Figure 4, label token probability and word intrusion confidence consistently yield better performance in terms of PN. This suggests that by including proposed confidence during topic refinement, we reduce the impact of potential noisy suggestions from the LLM and achieve better topical representation for documents. For further studies on alternative LLM confidence measures, see Appendix F.